OpenStudy (anonymous):

The length of a rectangular boat storage shed must be 4 meters more than the width, and perimeter must be at least 124 meters. What is the range of values of the width?

OpenStudy (anonymous):

can someone help out here

OpenStudy (anonymous):

Are you helping with this problem

jhonyy9 (jhonyy9):

perimeter of rectangle = 2l +2w l= w+4 so than P = 2(w+4)+2w = 2w+8+2w = 4w+8 but p need being least than 124 P<124 4w+8<124 4w<116 w<29 hope this will help you

OpenStudy (anonymous):

Thanks for your help